# Neural Approximate Mirror Maps for Constrained Diffusion Models

### v1

@@ A.3 Data assimilation with mirror DPS
for some unknown source image $\mathbf{x}^{*}\in\mathbb{R}^{d}$, where $\mathbf{\mathcal{A}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is a known forward operator, and $m\times m$ is the known noise covariance matrix $\mathbf{\Gamma}$. Higher values of $\zeta$ impose greater measurement consistency, but setting $\zeta$ too high can cause instabilities and artifacts. The data assimilation results in Fig. 5 used $\zeta=0.1$ and constraint-guidance weight equal to $200$. Fig. 9 shows results for the same tasks but different values of $\zeta$ in DPS and the constraint-guidance weight in CG-DPS.
@@
Appendix B Constraint details
@@
The total brightness, or flux, of a discrete image $\mathbf{x}\in\mathbb{R}^{d}$ is simply the sum of its pixel values: $V(\mathbf{x}):=\sum_{i=1}^{d}\mathbf{x}_{i}$. We use the constraint distance function
@@
where $\bar{V}\in\mathbb{R}_{\geq 0}$ is the target total brightness. The dataset used for this constraint contains images from general relativistic magneto-hydrodynamic (GRMHD) simulations [68] of SgrA* with a fixed field of view. The images (originally $400\times 400$) were resized to $64\times 64$ pixels and rescaled to have a total flux of $120$. The dataset consists of $100000$ training images and $100$ validation images.
@@
Burgers’ equation [7, 13] is a nonlinear PDE that is a useful model for fluid mechanics. We consider the equation for a viscous fluid $u=u(t,x)$ in one-dimensional space:
@@
where $u(0,x)$ is some initial condition $u_{0}(x)$, and $\nu\in\mathbb{R}_{\geq 0}$ is the viscosity coefficient. We use Crank-Nicolson [19] to discretize and approximately solve Eq. 10 by representing the solution on an $n_{x}\times n_{t}$ grid, where $n_{x}$ is the spatial discretization, and $n_{t}$ is the number of snapshots in time. Given an $n_{x}\times n_{t}$ image, we wish to verify that it could be a solution to Eq. 10 with the Crank-Nicolson discretization. Letting $\mathbf{x}\in\mathbb{R}^{n_{x}\times n_{t}}$ denote the 2D image, we formulate the following distance function for evaluating agreement with the Crank-Nicolson solver:
@@
where $f_{\text{C-N}}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ outputs the snapshot at the next time using Crank-Nicolson, and Pythonic notation is used for simplicity. Note that a finite-differences loss à la physics-informed neural networks (PINNs) [54] would also work, but then our data would have non-negligible constraint distances since Crank-Nicolson solutions do not strictly follow a low-order finite-differences approximation.
Using a Crank-Nicolson solver [19] implemented with Diffrax [41], we numerically solved the 1D Burgers’ equation (Eq. 10) with viscosity coefficient $\nu=0.5$. The initial conditions were sampled from a Gaussian process based on a Matérn kernel with smoothness parameter $1.5$ and length scale equal to $1.0$. We discretized the spatiotemporal domain into a $64\times 64$ grid covering the spatial extent $x\in[0,10]$ and time interval $t\in[0,8]$. We ran Crank-Nicolson with a time step of $\Delta t=0.025$ and saved every fifth step for a total of $64$ snapshots. We followed this process to create our 1D Burgers’ dataset of $10000$ training images and $1000$ validation images.
@@
where $\mathbf{u}=\mathbf{u}(x,y,t)$ is the 2D velocity field at spatial location $(x,y)$ and time $t$, $Re$ is the Reynolds number, $\rho$ is the density, $p$ is the pressure field, and $\mathbf{f}$ is the external forcing. Following Kochkov et al. [44] and Rozet and Louppe [56], we set $Re=10^{3}$, $\rho=1$, and $\mathbf{f}$ corresponding to Kolmogorov forcing [14, 10] with linear damping. We consider the spatial domain $[0,2\pi]^{2}$ with periodic boundary conditions and discretize it into a $64\times 64$ uniform grid. We used jax-cfd [44] to randomly sample divergence-free, spectrally filtered initial conditions and then solve the Navier-Stokes equations with the forward Euler integration method with $\Delta t=0.01$ time units. We saved a snapshot every $20$ time units for a total of $8$ snapshots in the time interval $[3,4.6]$. We represent the solution as a two-channel $128\times 256$ image showing the snapshots in left-to-right order. In total, the dataset consists of $10000$ training images and $1000$ validation images.
@@
which compares each pair of tiles, where $\mathbf{t}_{i}(\mathbf{x})$ denotes the $i$-th tile in the image. For our experiments, we consider $32\times 32$ unit cells that are tiled in a $2\times 2$ pattern to create $64\times 64$ images. Using the unit-cell generation code of Ogren et al. [52], we created a dataset of $30000$ training images and $300$ validation images.
@@
We demonstrate this constraint with astronomical images that contain a certain number of galaxies. In particular, we simulated $128\times 128$ images of radio galaxies with background noise [18], each of which has exactly eight ($\bar{c}=8$) galaxies with an SNR $\geq 15$ dB. The dataset consists of $10000$ training images and $1000$ validation images.
To train the counting CNN, we created a mixed dataset with images of $6,7,8,9,$ or $10$ galaxies that includes $10000$ training images and $1000$ validation images for each of the five labels. The CNN architecture was adapted from a simple MNIST classifier [1] with two convolutional layers followed by two dense layers with ReLU activations. The CNN was trained to minimize the mean squared error between the real-valued estimated count and the ground-truth count.

### v2

@@ D.3 Data assimilation with mirror DPS
for some unknown source image $\mathbf{x}^{*}\in\mathbb{R}^{d}$, where $\mathbf{\mathcal{A}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ is a known forward operator, and $m\times m$ is the known noise covariance matrix $\mathbf{\Gamma}$. Higher values of $\zeta$ impose greater measurement consistency, but setting $\zeta$ too high can cause instabilities and artifacts. The data assimilation results in Figure 5 used $\zeta=0.1$ and constraint-guidance weight equal to $200$. Figure 11 shows results for the same tasks but different values of $\zeta$ in DPS and the constraint-guidance weight in CG-DPS.
@@
Appendix E Measures of distance between distributions
E.1 MMD
The maximum mean discrepancy (MMD) between two distributions is computed by embedding both distributions into a reproducing kernel Hilbert space (RKHS) and using samples to estimate the resulting distance. We use the popular Gaussian radial basis function (RBF) kernel to construct the RKHS, setting the length scale $\sigma$ as
$\sqrt{\text{median}\left(\left\{\left\lVert\mathbf{x}^{(i)}-\mathbf{x}^{(j)}\right\rVert_{2}^{2}\right\}\right)/2},$ (10)
i.e., the square root of half the median of the pairwise squared Euclidean distances in the dataset $\left\{\mathbf{x}^{(i)}\right\}_{i=1}^{n}$. This is a popular choice in previous work (Briol et al., 2019) and has been theoretically and empirically justified Garreau et al. (2017).
Our MMD implementation is based on the code provided for the work of Sutherland et al. (2016) (Sutherland, 2018). We estimated mean and standard deviation empirically with $50$ random subsets of $1000$ samples from each dataset. The length scale was estimated (Equation 10) for each subset using the samples in the true subset. In total, the generated and true datasets contained $10000$ samples each. Held-out test images were used as true samples for all constraints, except for total brightness (due to a lack of test images in the dataset, training images were used for this constraint only).
E.2 KID
The Kernel Inception Distance (KID) between two distributions is based on the Inception v3 features evaluated for samples from both distributions. Following standard practice, we use the $2048$-dimensional final average pooling features. The KID is computed as the squared MMD (using a polynomial kernel) between the two embedded distributions. It has several advantages over the Fréchet Inception Distance (FID) (Heusel et al., 2017), including being unbiased and more sample-efficient (Bińkowski et al., 2018). We note, however, that the Inception network was trained on natural images, so both KID and FID are not perfect metrics for the types of data we consider in this work, such as physics-based simulation outputs.
KID evaluation is based on the gan-metrics-pytorch repository (Fatir, 2021), using the same Inception v3 weights as those used in the official TensorFlow implementation of FID (Heusel et al., 2017). We evaluated KID with the same samples that were used for MMD. Since the Inception network takes RGB images as input, we represented the samples as grayscale images converted to RGB. For all the constraint datasets except for the divergence-free Kolmogorov flows, we clipped the image pixel values to $[0,1]$ before converting them to RGB. For the divergence-free data, we clipped the values in the vorticity images to $[-20,20]$ and then rescaled this range to $[0,1]$.
E.3 FID
Although FID is a biased finite-sample estimator and heavily depends on the number of samples, it is a popular metric for evaluating generative models. Table 5 reports FID values in addition to the MMD and KID values in Table 1. We find that the rankings of our method before finetuning, our method after finetuning, and the baseline method are consistent with the rankings when using KID in Table 1. Again we note that FID and KID are based on Inception features that were tuned to natural images, so they are not the most reliable measures of distance between distributions for these particular image datasets.
Total Brightness 1D Burgers’ Divergence-free Periodic Count
Ours w/o FT $5.02$ $3.36$ $\mathbf{2.21}$ $2.26$ $\mathbf{24.18}$
Ours $\mathbf{4.58}$ $\mathbf{2.04}$ $3.97$ $\mathbf{1.99}$ $37.65$
Baseline $42.25$ $140.70$ $3.41$ $6.42$ $69.76$
Table 5: FID values ($\downarrow$) associated with the comparisons in Tab. 1. According to FID, our approach consistently outperforms the baseline DM approach in matching the true distribution. Finetuning even sometimes improves FID while improving constraint satisfaction (see Tab. 1 for constraint distances).
Our FID implementation is borrowed from the pytorch-fid codebase (Seitzer, 2020). Per standard practice, we estimated FID with $50000$ generated samples and $50000$ true samples. As was the case with MMD and KID, held-out test images were used as true samples, except that training images were used for the total brightness data. We pre-processed the samples for the Inception network in the same way we did for KID.
Appendix F Constraint details
@@
The total brightness, or total flux, of a discrete image $\mathbf{x}\in\mathbb{R}^{d}$ is simply the sum of its pixel values: $V(\mathbf{x}):=\sum_{i=1}^{d}\mathbf{x}_{i}$. We use the constraint distance function
@@
where $\bar{V}\in\mathbb{R}_{\geq 0}$ is the target total brightness. The dataset used for this constraint contains images from general relativistic magneto-hydrodynamic (GRMHD) simulations (Wong et al., 2022) of Sgr A* with a fixed field of view. The images (originally $400\times 400$) were resized to $64\times 64$ pixels and rescaled to have a total flux of $120$. The dataset consists of $100000$ training images and $100$ validation images.
@@
Burgers’ equation (Bateman, 1915; Burgers, 1948) is a nonlinear PDE that is a useful model for fluid mechanics. We consider the equation for a viscous fluid $u=u(t,x)$ in one-dimensional space:
@@
where $u(0,x)$ is some initial condition $u_{0}(x)$, and $\nu\in\mathbb{R}_{\geq 0}$ is the viscosity coefficient. We use Crank-Nicolson (Crank & Nicolson, 1947) to discretize and approximately solve Equation 11 by representing the solution on an $n_{x}\times n_{t}$ grid, where $n_{x}$ is the spatial discretization, and $n_{t}$ is the number of snapshots in time. Given an $n_{x}\times n_{t}$ image, we wish to verify that it could be a solution to Equation 11 with the Crank-Nicolson discretization. Letting $\mathbf{x}\in\mathbb{R}^{n_{x}\times n_{t}}$ denote the 2D image, we formulate the following distance function for evaluating agreement with the Crank-Nicolson solver:
@@
where $f_{\text{C-N}}:\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$ outputs the snapshot at the next time using Crank-Nicolson, and Pythonic notation is used for simplicity. Note that a finite-differences loss as proposed for physics-informed neural networks (PINNs) Raissi et al. (2019) would also work, but then our data would have non-negligible constraint distances since Crank-Nicolson solutions do not strictly follow a low-order finite-differences approximation.
Using a Crank-Nicolson solver (Crank & Nicolson, 1947) implemented with Diffrax (Kidger, 2022), we numerically solved the 1D Burgers’ equation (Equation 11) with viscosity coefficient $\nu=0.5$. The initial conditions were sampled from a Gaussian process based on a Matérn kernel with smoothness parameter $1.5$ and length scale equal to $1.0$. We discretized the spatiotemporal domain into a $64\times 64$ grid covering the spatial extent $x\in[0,10]$ and time interval $t\in[0,8]$. We ran Crank-Nicolson with a time step of $\Delta t=0.025$ and saved every fifth step for a total of $64$ snapshots. We followed this process to create our 1D Burgers’ dataset of $10000$ training images and $1000$ validation images.
@@
where $\mathbf{u}=\mathbf{u}(x,y,t)$ is the 2D velocity field at spatial location $(x,y)$ and time $t$, $Re$ is the Reynolds number, $\rho$ is the density, $p$ is the pressure field, and $\mathbf{f}$ is the external forcing. Following Kochkov et al. (2021) and Rozet & Louppe (2024), we set $Re=10^{3}$, $\rho=1$, and $\mathbf{f}$ corresponding to Kolmogorov forcing (Chandler & Kerswell, 2013; Boffetta & Ecke, 2012) with linear damping. We consider the spatial domain $[0,2\pi]^{2}$ with periodic boundary conditions and discretize it into a $64\times 64$ uniform grid. We used jax-cfd (Kochkov et al., 2021) to randomly sample divergence-free, spectrally filtered initial conditions and then solve the Navier-Stokes equations with the forward Euler integration method with $\Delta t=0.01$ time units. We saved a snapshot every $20$ time units for a total of $8$ snapshots in the time interval $[3,4.6]$. We represent the solution as a two-channel $128\times 256$ image showing the snapshots in left-to-right order. In total, the dataset consists of $10000$ training images and $1000$ validation images.
@@
which compares each pair of tiles, where $\mathbf{t}_{i}(\mathbf{x})$ denotes the $i$-th tile in the image. For our experiments, we consider $32\times 32$ unit cells that are tiled in a $2\times 2$ pattern to create $64\times 64$ images. Using the unit-cell generation code of Ogren et al. (2024), we created a dataset of $30000$ training images and $300$ validation images.
@@
We demonstrate this constraint with astronomical images that contain a certain number of galaxies. In particular, we simulated $128\times 128$ images of radio galaxies with background noise Connor et al. (2022), each of which has exactly eight ($\bar{c}=8$) galaxies with an SNR $\geq 15$ dB. The dataset consists of $10000$ training images and $1000$ validation images.
To train the counting CNN, we created a mixed dataset with images of $6,7,8,9,$ or $10$ galaxies that includes $10000$ training images and $1000$ validation images for each of the five labels. The CNN architecture was adapted from a simple MNIST classifier (8bitmp3, 2023) with two convolutional layers followed by two dense layers with ReLU activations. The CNN was trained to minimize the mean squared error between the real-valued estimated count and the ground-truth count.